### v1

# Mathematical proof about spatial symmetry of solutions of the two-dimensional Kolmogorov flow

Abstract
We give a mathematical proof that solution for all $t>0$ of the two-dimensional (2D) Kolmogorov flow governed by Navier-Stokes (NS) equations with periodic boundary condition remains the same spatial symmetry as its smooth initial condition at $t=0$. This mathematical theorem supports the corresponding CNS (clean numerical simulation) results of the 2D turbulent Kolmogorov flow [20, 8] that remain the same spatial symmetry, but does not support the corresponding DNS (direct numerical simulation) results that lose the spatial symmetry quickly. In other words, these DNS results violate this mathematical theorem. Thus, this mathematical theorem rigorously confirms that the spatiotemporal trajectories of NS turbulence given by DNS are indeed quickly polluted by numerical noises badly. It also illustrates that CNS can provide helpful enlightenments to deepen our understanding about turbulence and besides approach some mathematical truths about NS equations.
Keyword Navier-Stokes equations, spatial symmetry, turbulence, CNS, DNS
1 Introduction
Lorenz, who put forward in 1963 the famous “butterfly-effect” of chaos [14], further discovered in 2006 that numerical errors could have significant influences on global properties of chaotic systems [15], because unavoidable numerical noises as small artificial disturbance exponentially increase to macro-level. Note that the Navier-Stokes turbulence (i.e. turbulence governed by NS equations) is chaotic [5, 6, 2, 1, 3]. However, for all traditional numerical algorithms including the direct numerical simulation (DNS) of NS equations, numerical noises are unavoidable. Thus, logically speaking, numerical noises of DNS of NS turbulence should quickly increase to a macro-level, which certainly should lead to a distinct departure from exact trajectory of NS equations. Thus, DNS of NS turbulence should be quickly polluted by numerical noises badly.
Liao [12] proposed the so-called “clean numerical simulation” (CNS) to overcome the restrictions of the traditional numerical algorithms for chaos and turbulence. Unlike DNS, numerical noises of CNS can be rigorously negligible in a finite time-interval, which is however long enough for statistics, by means of reducing both truncation error and round-off error to a required small enough level [13, 7, 10, 9, 17, 20, 8, 18]. Thus, CNS can give very accurate spatiotemporal trajectory of NS equations in a finite but long enough interval of time: they are very close to the exact solution.
Note that the numerical noises of CNS are much smaller than those of DNS. So, CNS results can be used as benchmark solution to check the reliability of DNS results. In fact, some comparisons were given for two-dimensional (2D) incompressible Kolmogorov flow [20, 8]: the numerical simulations given by CNS remain the same spatial symmetry (in a finite but long enough time-interval) as the corresponding initial condition, but the corresponding DNS results remain the same spatial symmetry only at the beginning but quickly lose it. Thus, spatial symmetry becomes a key point to check the reliability of numerical simulation of NS turbulence.
In this paper, we give mathematical proofs that solution for all $t>0$ of the 2D Kolmogorov flow considered in [20, 8] remains the same spatial symmetry as its initial condition. Note that the CNS results in [20, 8] agree well with this proof, but the corresponding DNS results lose the spatial symmetry and thus violate this mathematical theorem, clearly confirming that DNS results are indeed quickly polluted by numerical noises badly. The detailed mathematical proofs are given in § 2, the concluding remarks and discussions are given in § 3.
2 Spatial symmetry of 2D Kolmogorov flow
Consider a 2D incompressible Kolmogorov flow [16, 4, 21] in a square domain under the so-called Kolmogorov forcing, governed by the dimensionless Navier-Stokes equation in the form of stream-function
$\frac{\partial}{\partial t}\Big(\nabla^{2}\psi\Big)+\psi_{x}\nabla^{2}\psi_{y}-\psi_{y}\nabla^{2}\psi_{x}-\frac{1}{Re}\nabla^{4}\psi+n_{K}\cos(n_{K}y)=0,$ (1)
where $\psi$ is the stream-function, $t$ denotes the time, $x,y\in[0,2\pi]$ are horizontal and vertical coordinates, $Re$ is the Reynolds number, $n_{K}$ is an even number describing the Kolmogorov forcing scale, $\nabla^{2}$ is the Laplace operator, and $\nabla^{4}=\nabla^{2}\nabla^{2}$, respectively. The stream-function $\psi$ satisfies the periodic boundary condition
$\psi(x,y,t)=\psi(x\pm 2\pi,y,t)=\psi(x,y\pm 2\pi,t).$ (2)
Here, let us consider an arbitrarily smooth initial condition $\psi(x,y,0)$ that contains the spatial symmetry of rotation
$\psi(x,y,0)=\psi(2\pi-x,2\pi-y,0)$ (3)
and/or translation
$\psi(x,y,0)=\psi(x+\pi,y+\pi,0).$ (4)
Assume that $\psi(x,y,t)$ can be expanded as Taylor series at $t=t_{0}\geq 0$ with a non-zero radius $\rho>0$ of convergence, say,
$\psi(x,y,t)$ $=$ $\psi(x,y,t_{0})+\sum_{m=1}^{+\infty}\psi^{(m)}(x,y,t_{0})(t-t_{0})^{m}$ (5)
$=$ $\sum_{m=0}^{+\infty}\psi^{(m)}(x,y,t_{0})(t-t_{0})^{m},$
where $|t-t_{0}|<\rho$, and
$\psi^{(m)}(x,y,t_{0})=\left.\frac{1}{m!}\frac{\partial^{m}\psi(x,y,t)}{\partial t^{m}}\right|_{t=t_{0}}$ (6)
with the definition
$\psi^{(0)}(x,y,t_{0})=\psi(x,y,t_{0}).$ (7)
Then, we can rewrite (1) as follows
$\nabla^{2}\psi^{(1)}(x,y,t)+\psi^{(0)}_{x}\nabla^{2}\psi^{(0)}_{y}-\psi^{(0)}_{y}\nabla^{2}\psi^{(0)}_{x}-\frac{1}{Re}\nabla^{4}\psi^{(0)}+n_{K}\cos(n_{K}y)=0.$ (8)
Under the coordinate transformation
$x^{\prime}=2\pi-x,\;\;y^{\prime}=2\pi-y$ (9)
for the spatial symmetry of rotation, and
$x^{\prime}=\pi+x,\;\;y^{\prime}=\pi+y$ (10)
for the spatial symmetry of translation, respectively, the corresponding solution reads $\psi^{\prime}(x^{\prime},y^{\prime},t)$. Since there exists the spatial symmetry of rotation (3) and/or translation (4), it holds
$\psi^{\prime}(x^{\prime},y^{\prime},0)=\psi(x,y,0).$ (11)
theorem 1
For the two-dimensional Kolmogorov flow governed by the Navier-Stokes equation (1) (with an even number $n_{K}$) under the periodic boundary condition (2) subject to a smooth initial condition with the spatial symmetry of rotation (3) and/or translation (4), if its solution $\psi(x,y,t)$ at $t=t_{0}\geq 0$ contains a spatial symmetry $\psi(x,y,t_{0})=\psi(x^{\prime},y^{\prime},t_{0})$ of rotation and/or translation, then there exists the spatial symmetry $\psi^{(m)}(x,y,t_{0})=\psi^{(m)}(x^{\prime},y^{\prime},t_{0})$ for all integers $m\geq 1$.
Proof: In the case of (9), we have
$\frac{\partial}{\partial x}=\frac{\partial}{\partial x^{\prime}}\frac{\partial x^{\prime}}{\partial x}=-\frac{\partial}{\partial x^{\prime}},\;\;\;\frac{\partial}{\partial y}=\frac{\partial}{\partial y^{\prime}}\frac{\partial y^{\prime}}{\partial y}=-\frac{\partial}{\partial y^{\prime}},$ (12)
$\frac{\partial^{2}}{\partial x^{2}}=-\frac{\partial}{\partial x^{\prime}}\left(-\frac{\partial}{\partial x^{\prime}}\right)=\frac{\partial^{2}}{\partial x^{\prime 2}},\;\;\frac{\partial^{2}}{\partial y^{2}}=-\frac{\partial}{\partial y^{\prime}}\left(-\frac{\partial}{\partial y^{\prime}}\right)=\frac{\partial^{2}}{\partial y^{\prime 2}},$ (13)
so that
$\nabla^{2}$ $=$ $\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}=\frac{\partial^{2}}{\partial x^{\prime 2}}+\frac{\partial^{2}}{\partial y^{\prime 2}}=\nabla^{\prime 2},$ (14)
$\nabla^{4}=\nabla^{2}\;\nabla^{2}=\nabla^{\prime 2}\;\nabla^{\prime 2}=\nabla^{\prime 4},$ (15)
and
$\frac{\partial f}{\partial x}\;\frac{\partial g}{\partial y}=\left(-\frac{\partial f}{\partial x^{\prime}}\right)\left(-\frac{\partial g}{\partial y^{\prime}}\right)=\frac{\partial f}{\partial x^{\prime}}\;\frac{\partial g}{\partial y^{\prime}}$ (16)
for smooth functions $f$ and $g$.
Similarly, in the case of (10), we have
$\frac{\partial}{\partial x}=\frac{\partial}{\partial x^{\prime}},\;\;\;\frac{\partial}{\partial y}=\frac{\partial}{\partial y^{\prime}},\;\;\;\frac{\partial^{2}}{\partial x^{2}}=\frac{\partial^{2}}{\partial x^{\prime 2}},\;\;\;\frac{\partial^{2}}{\partial y^{2}}=\frac{\partial^{2}}{\partial y^{\prime 2}},$ (17)
so that
$\nabla^{2}=\nabla^{\prime 2},\nabla^{4}=\nabla^{\prime 4}$ (18)
and
$\frac{\partial f}{\partial x}\;\frac{\partial g}{\partial y}=\frac{\partial f}{\partial x^{\prime}}\;\frac{\partial g}{\partial y^{\prime}}$ (19)
for smooth functions $f$ and $g$.
Besides, in the case of (9), it holds
$\cos(n_{K}y^{\prime})=\cos(2n_{K}\pi-n_{k}y)=\cos(n_{K}y).$ (20)
In the case of (10), it holds
$\cos(n_{K}y^{\prime})=\cos(n_{K}\pi+n_{k}y)=\cos(n_{K}y),$ (21)
since $n_{k}$ is an even number.
Thus, in the case of (9) and/or (10), it holds
$\cos(n_{K}y)=\cos(n_{K}y^{\prime}),\;\;\nabla^{2}=\nabla^{\prime 2},\;\;\nabla^{4}=\nabla^{\prime 4},\;\;\frac{\partial f}{\partial x}\;\frac{\partial g}{\partial y}=\frac{\partial f}{\partial x^{\prime}}\;\frac{\partial g}{\partial y^{\prime}}$ (22)
for smooth functions $f$ and $g$.
(A) When $m=1$
Since there exists the spatial symmetry at $t=t_{0}$, it holds
$\psi^{\prime}(x^{\prime},y^{\prime},t_{0})=\psi(x,y,t_{0}),\mbox{i.e.}\hskip 8.5359pt\psi^{\prime(0)}(x^{\prime},y^{\prime},t_{0})=\psi^{(0)}(x,y,t_{0}).$ (23)
Since $\psi^{\prime}(x^{\prime},y^{\prime},t_{0})$ satisfies the same equations, we have
$\nabla^{\prime 2}\psi^{\prime(1)}(x^{\prime},y^{\prime},t)+\psi^{\prime(0)}_{x^{\prime}}\nabla^{\prime 2}\psi^{\prime(0)}_{y^{\prime}}-\psi^{\prime(0)}_{y^{\prime}}\nabla^{\prime 2}\psi^{\prime(0)}_{x^{\prime}}-\frac{1}{Re}\nabla^{\prime 4}\psi^{\prime(0)}+n_{K}\cos(n_{K}y^{\prime})=0.$ (24)
According to (22) and (23), we have at $t=t_{0}$ that
$\nabla^{\prime 2}\psi^{\prime(1)}(x^{\prime},y^{\prime},t)=\nabla^{2}\psi^{\prime(1)}(x^{\prime},y^{\prime},t),\;\;\;\cos(n_{K}y^{\prime})=\cos(n_{K}y)$ (25)
and
$\psi^{\prime(0)}_{x^{\prime}}\nabla^{\prime 2}\psi^{\prime(0)}_{y^{\prime}}-\psi^{\prime(0)}_{y^{\prime}}\nabla^{\prime 2}\psi^{\prime(0)}_{x^{\prime}}-\frac{1}{Re}\nabla^{\prime 4}\psi^{\prime(0)}$ (26)
$=$ $\psi^{\prime(0)}_{x^{\prime}}\left(\nabla^{\prime 2}\psi^{\prime(0)}\right)_{y^{\prime}}-\psi^{\prime(0)}_{y^{\prime}}\left(\nabla^{\prime 2}\psi^{\prime(0)}\right)_{x^{\prime}}-\frac{1}{Re}\nabla^{4}\psi^{\prime(0)}$
$=$ $\psi^{\prime(0)}_{x^{\prime}}\left(\nabla^{2}\psi^{\prime(0)}\right)_{y^{\prime}}-\psi^{\prime(0)}_{y^{\prime}}\left(\nabla^{2}\psi^{\prime(0)}\right)_{x^{\prime}}-\frac{1}{Re}\nabla^{4}\psi^{\prime(0)}$
$=$ $\psi^{(0)}_{x}\left(\nabla^{2}\psi^{(0)}\right)_{y}-\psi^{(0)}_{y}\left(\nabla^{2}\psi^{(0)}\right)_{x}-\frac{1}{Re}\nabla^{4}\psi^{(0)}$
$=$ $\psi^{(0)}_{x}\nabla^{2}\psi^{(0)}_{y}-\psi^{(0)}_{y}\nabla^{2}\psi^{(0)}_{x}-\frac{1}{Re}\nabla^{4}\psi^{(0)},$
so that (24) becomes
$\nabla^{2}\psi^{\prime(1)}(x^{\prime},y^{\prime},t)+\psi^{(0)}_{x}\nabla^{2}\psi^{(0)}_{y}-\psi^{(0)}_{y}\nabla^{2}\psi^{(0)}_{x}-\frac{1}{Re}\nabla^{4}\psi^{(0)}+n_{K}\cos(n_{K}y)=0.$ (27)
Comparing (27) to (8) gives the spatial symmetry
$\psi^{\prime(1)}(x^{\prime},y^{\prime},t_{0})=\psi^{(1)}(x,y,t_{0}).$ (28)
(B) For a given positive integer $m$
Assume that for a given positive integer $m$ there exists the spatial symmetry at $t=t_{0}$, say,
$\psi^{\prime(i)}(x^{\prime},y^{\prime},t_{0})=\psi^{(i)}(x,y,t_{0}),\hskip 28.45274pt0\leq i\leq m.$ (29)
Differentiating (1) and (24) $m$ times with respect to $t$ and then dividing by $m!$, we have
$(m+1)\nabla^{2}\psi^{(m+1)}(x,y,t)$ (30)
$=$ $\frac{1}{Re}\nabla^{4}\psi^{(m)}(x,y,t)-\sum_{i=0}^{m}\Big[\psi^{(i)}_{x}\nabla^{2}\psi^{(m-i)}_{y}-\psi^{(i)}_{y}\nabla^{2}\psi^{(m-i)}_{x}\Big$]
and
$(m+1)\nabla^{\prime 2}\psi^{\prime(m+1)}(x^{\prime},y^{\prime},t)$ (31)
$=$ $\frac{1}{Re}\nabla^{\prime 4}\psi^{\prime(m)}(x^{\prime},y^{\prime},t)-\sum_{i=0}^{m}\left[\psi^{\prime(i)}_{x^{\prime}}\nabla^{\prime 2}\psi^{\prime(m-i)}_{y^{\prime}}-\psi^{\prime(i)}_{y^{\prime}}\nabla^{\prime 2}\psi^{\prime(m-i)}_{x^{\prime}}\right$],
respectively.
According to (22) and (29), we have at $t=t_{0}$ that
$\nabla^{\prime 2}\psi^{\prime(m+1)}=\nabla^{2}\psi^{\prime(m+1)},\;\;\;\nabla^{\prime 4}\psi^{\prime(m)}=\nabla^{4}\psi^{(m)},$ (32)
and
$\sum_{i=0}^{m}\left[\psi^{\prime(i)}_{x^{\prime}}\nabla^{\prime 2}\psi^{\prime(m-i)}_{y^{\prime}}-\psi^{\prime(i)}_{y^{\prime}}\nabla^{\prime 2}\psi^{\prime(m-i)}_{x^{\prime}}\right$] (33)
$=$ $\sum_{i=0}^{m}\left[\psi^{\prime(i)}_{x^{\prime}}\left(\nabla^{\prime 2}\psi^{\prime(m-i)}\right)_{y^{\prime}}-\psi^{\prime(i)}_{y^{\prime}}\left(\nabla^{\prime 2}\psi^{\prime(m-i)}\right)_{x^{\prime}}\right]$
$=$ $\sum_{i=0}^{m}\left[\psi^{\prime(i)}_{x^{\prime}}\left(\nabla^{2}\psi^{\prime(m-i)}\right)_{y^{\prime}}-\psi^{\prime(i)}_{y^{\prime}}\left(\nabla^{2}\psi^{\prime(m-i)}\right)_{x^{\prime}}\right]$
$=$ $\sum_{i=0}^{m}\left[\psi^{(i)}_{x}\left(\nabla^{2}\psi^{(m-i)}\right)_{y}-\psi^{(i)}_{y}\left(\nabla^{2}\psi^{(m-i)}\right)_{x}\right]$
$=$ $\sum_{i=0}^{m}\Big[\psi^{(i)}_{x}\nabla^{2}\psi^{(m-i)}_{y}-\psi^{(i)}_{y}\nabla^{2}\psi^{(m-i)}_{x}\Big$].
Substituting (32) and (33) into (31) gives at $t=t_{0}$ that
$(m+1)\nabla^{2}\psi^{\prime(m+1)}(x^{\prime},y^{\prime},t_{0})$ (34)
$=$ $\frac{1}{Re}\nabla^{4}\psi^{(m)}(x,y,t_{0})-\sum_{i=0}^{m}\Big[\psi^{(i)}_{x}\nabla^{2}\psi^{(m-i)}_{y}-\psi^{(i)}_{y}\nabla^{2}\psi^{(m-i)}_{x}\Big$].
Comparing (34) to (30) gives the spatial symmetry
$\psi^{\prime(m+1)}(x^{\prime},y^{\prime},t_{0})=\psi^{(m+1)}(x,y,t_{0})$ (35)
for a given positive integer $m$.
(C) For an arbitrary integer $m\geq 0$
According to (A) and (B), it holds the spatial symmetry
$\psi^{\prime(m)}(x^{\prime},y^{\prime},t_{0})=\psi^{(m)}(x,y,t_{0}),$ (36)
for an arbitrary integer $m\geq 0$. This ends the proof of theorem 1.
theorem 2
For the two-dimensional Kolmogorov flow governed by the Navier-Stokes equation (1) (with an even number $n_{K}$) under the periodic boundary condition (2) subject to a smooth initial condition with the spatial symmetry of rotation (3) and/or translation (4), if its solution $\psi(x,y,t)$ at $t=t_{0}$ contains a spatial symmetry of rotation and/or translation, say, $\psi^{\prime}(x^{\prime},y^{\prime},t_{0})=\psi(x,y,t_{0})$, and besides the radius $\rho$ of convergence of its Taylor series at $t=t_{0}$ is non-zero, then $\psi(x,y,t)$ in $t\in[t_{0},t_{0}+\rho)$ has the same spatial symmetry as $\psi(x,y,t_{0})$.
Proof: According to theorem 1, from the spatial symmetry $\psi^{\prime}(x^{\prime},y^{\prime},t_{0})=\psi(x,y,t_{0})$, we have the spatial symmetry
$\psi^{\prime(m)}(x^{\prime},y^{\prime},t_{0})=\psi^{(m)}(x,y,t_{0}),\hskip 28.45274pt1\leq m<+\infty.$ (37)
Then,
$\psi^{\prime}(x^{\prime},y^{\prime},t)$ (38)
$=$ $\sum_{m=0}^{+\infty}\psi^{\prime(m)}(x^{\prime},y^{\prime},t_{0})\;(t-t_{0})^{m}=\sum_{m=0}^{+\infty}\psi^{(m)}(x,y,t_{0})(t-t_{0})^{m}$
$=$ $\psi(x,y,t),$
where $|t-t_{0}|<\rho$, including $t\in[t_{0},t_{0}+\rho)$, where $\rho>0$ is the radius of convergence. This guarantees the spatial symmetry $\psi^{\prime}(x^{\prime},y^{\prime},t)=\psi(x,y,t)$ in $t\in[t_{0},t_{0}+\rho)$. This ends the proof of theorem 2.
theorem 3
For the two-dimensional Kolmogorov flow governed by the Navier-Stokes equation (1) (with an even number $n_{K}$) under the periodic boundary condition (2) subject to a smooth initial condition with the spatial symmetry of rotation (3) and/or translation (4), if the convergence radius $\rho$ of its Taylor series is non-zero for all $t\geq 0$, its solution for all $t>0$ has the same spatial symmetry as its initial condition.
Proof: Since the initial condition contains the spatial symmetry, i.e. $\psi(x,y,0)=\psi^{\prime}(x^{\prime},y^{\prime},0)$, we have according to theorem 2 (by setting $t_{0}=0$) the spatial symmetry
$\psi(x,y,t)=\psi^{\prime}(x^{\prime},y^{\prime},t),\hskip 28.45274ptt\in[0,\delta_{0}],$
where $\delta_{0}<\rho_{0}$, and $\rho_{0}$ is the radius of convergence of its Taylor series at $t=0$. Then, setting $t_{0}=\delta_{0}$, we have according to theorem 2 the spatial symmetry
$\psi(x,y,t)=\psi^{\prime}(x^{\prime},y^{\prime},t),\hskip 28.45274ptt\in[\delta_{0},\delta_{1}],$
where $\delta_{1}<\rho_{1}$, and $\rho_{1}$ is the radius of convergence of its Taylor series at $t=t_{0}=\delta_{0}$. Furthermore, setting $t_{0}=\delta_{1}$, we have according to theorem 2 the spatial symmetry
$\psi(x,y,t)=\psi^{\prime}(x^{\prime},y^{\prime},t),\hskip 28.45274ptt\in[\delta_{1},\delta_{2}],$
where $\delta_{2}<\rho_{2}$, and $\rho_{2}$ is the radius of convergence of its Taylor series at $t=t_{0}=\delta_{1}$. Note that
$0,\delta_{0}]\cup[\delta_{0},\delta_{1}]\cup[\delta_{1},\delta_{2}]=[0,\delta_{2}],$
thus now there exists the spatial symmetry
$\psi(x,y,t)=\psi^{\prime}(x^{\prime},y^{\prime},t),\hskip 28.45274ptt\in[0,\delta_{2}].$
The same process can keep going, since the Taylor series of $\psi(x,y,t)$ always exists with a non-zero radius of convergence. Thus, it holds the spatial symmetry
$\psi(x,y,t)=\psi^{\prime}(x^{\prime},y^{\prime},t),\;\;t\in[0,+\infty).$
This ends the proof of the theorem 3.
3 Concluding remarks and discussions
As reported in [20, 8], the numerical simulations of the 2D Kolmogorov flow given by CNS (in a finite but long enough time-interval) always remain the same spatial symmetry as the corresponding initial condition. This agrees well with the theorem 3. This is mainly because, unlike traditional algorithms, numerical noises of CNS are rigorously negligible in a finite time-interval that is however long enough for statistics.
However, as mentioned in [20, 8], the corresponding DNS results remain the same spatial symmetry as the corresponding initial condition only in a short duration from the beginning but then quickly lose the spatial symmetry, clearly indicating that DNS results distinctly deviate the exact solution of the NS equations. In other words, these DNS results violate the theorem 3. This clearly indicates that the DNS results are indeed quickly polluted by numerical noises badly. The above conclusion should have general meanings for turbulent flows, as illustrated in [17].
Liao and Qin [8] applied CNS to solve the 2D turbulent Kolmogorov flow (1) in the case of $Re=2000$ and $n_{K}=16$, subject to the different initial conditions
$\psi_{1}(x,y,0)$ $=$ $-\frac{1}{2}\left[\cos(x+y)+\cos(x-y)\right],$ (39)
$\psi_{2}(x,y,0)$ $=$ $-\frac{1}{2}\left[\cos(x+y)+\cos(x-y)\right]+\delta^{\prime}\;\sin(x+y),$ (40)
where $\delta^{\prime}$ is a constant. Note that $\psi_{1}(x,y,0)$ contains the spatial symmetry of both rotation and translation, but $\psi_{2}(x,y,0)$ contains the spatial symmetry of only translation for a large $\delta^{\prime}\sim O(1)$. As reported by Liao and Qin [8], the spatiotemporal trajectories given by CNS using the initial condition $\psi_{1}(x,y,0)$ remains the same spatial symmetry of rotation and translation as $\psi_{1}(x,y,0)$ in the whole time-interval $t\in[0,300]$. This agrees well with the theorem 3 proved in this paper. When $\delta^{\prime}=10^{{-20}}$ in (40), the spatiotemporal trajectories given by CNS using the initial condition $\psi_{2}(x,y,0)$ remains the same spatial symmetry of rotation and translation as the initial condition $\psi_{1}(x,y,0)$ at the beginning, mainly because the disturbance $10^{-20}\sin(x+y)$ of $\psi_{2}(x,y,0)$ is too small compared to $\psi_{1}(x,y,0)$, until it enlarges to a macro-level, due to the so-called “butterfly-effect” of the chaotic characteristics of turbulence [5, 6, 2, 1, 3], at about $t\approx 35$ when it finally destroys the original spatial symmetry of $\psi_{1}(x,y,0)$, thus the CNS result thereafter remains the same spatial symmetry of translation as its initial condition $\psi_{2}(x,y,0)$ for large $\delta^{\prime}\sim O(1)$, as reported by Liao and Qin [8]. This CNS result also agrees well with the theorem 3. Certainly, the two CNS results should have different statistics, since they have different spatial symmetries. So, the theorem 3 also supports the conclusion that the turbulence is a kind of ultra-chaos [10, 19], i.e. small disturbance could lead to distinct differences not only in spatotemporal trajectory but also in statistics of turbulence. It should be emphasized again that, the corresponding DNS results of the same turbulent Kolmogorov flows quickly lose their spatial symmetry, since random numerical noises without spatial symmetry quickly enlarges to a macro-level. Therefore, the theorem 3 rigorously confirms that the DNS results of the 2D turbulent Kolmogorov flow (1) are indeed quickly polluted by numerical noises badly, as illustrated in [20]. This is also the reason why the so-called “noise-expansion cascade” revealed in [8] by means of CNS cannot be discovered by means of DNS.
For a large enough Reynolds number such as $Re=2000$ considered in [8], the 2D Kolmogorov flow (1) is turbulent. It is widely accepted by fluid community that turbulence is chaotic [5, 6, 2, 1, 3] and thus, due to the famous butterfly-effect of chaotic system, the small disturbance $\delta^{\prime}\sin(x+y)$ of the initial condition $\psi_{2}(x,y,t)$ defined by (40) should be quickly enlarged to the same order of magnitude as the first initial condition $\psi_{1}(x,y,t)$ described by (39). However, $\psi_{1}(x,y,t)$ has a different spatial symmetry from that of $\psi_{2}(x,y,t)$ when $\delta^{\prime}\sim O(1)$, therefore, according to the theorem 3, the turbulent Kolmogorov flow given by $\psi_{1}(x,y,t)$ has different spatial symmetry from that given by $\psi_{2}(x,y,t)$ even for arbitrarily small $\delta^{\prime}$, say, $\delta^{\prime}\to 0$. Thus, the theorem 3 plus the butterfly-effect of chaos supports the conjecture mentioned in [11]:
Conjecture A
The Navier-Stokes equations admit distinct smooth global solutions for all $t>0$ from almost the same initial conditions $U_{1}({\bf r})$ and $U_{2}({\bf r})$ with $|U_{1}({\bf r})-U_{2}({\bf r})|\leq\delta$ for arbitrarily small $\delta$ (or as $\delta\to 0$).
Note that the 2D Kolmogorov flow is a special case of NS equations. If one can give a mathematical proof about the butterfly-effect of the turbulence, say, the small disturbance $\delta^{\prime}\sin(x+y)$ of the 2D turbulent Kolmogorov flow subject to the initial condition $\psi_{2}(x,y,t)$ defined by (40) quickly enlarges to $O(1)$ for arbitrary small $\delta^{\prime}$, then, using the theorem 3, one can prove the following conjecture:
Conjecture B
The two-dimensional turbulent Kolmogorov flow admit distinct smooth global solutions for all $t>0$ from almost the same initial conditions $\psi_{1}(x,y,0)$ and $\psi_{2}(x,y,0)$ with $|\psi_{1}(x,y,0)-\psi_{2}(x,y,0)|\leq\delta$ for arbitrarily small $\delta$ (or as $\delta\to 0$).
Obviously, if the flow is laminar, all small disturbances in the initial condition should not increase to a macro-level, say, remain small forever. Thus, the above two conjectures should hold for some turbulent flows.
As reported by Qin and Liao [18], the CNS results of the three-dimensional (3D) Kolmogorov flow also remain the same spatial symmetry as its initial condition in a finite but long enough interval of time $t\in[0,500]$. Thus, it is quite possible that there should exist a corresponding theorem similar to theorem 3 for the 3D Kolmogorov flow, although it is an open question whether or not there should exist such a similar theorem for general NS turbulence under periodic boundary conditions.
Note that, unlike the theorem 3 proved in this paper, CNS results are reliable only in a finite but long enough interval of time $t\in[0,T_{c}]$, where $T_{c}$ is called the “critical predictable time”. As pointed out by Liao in his book [13],
$T_{c}\sim-\kappa\ln{\cal E}_{0},$ (41)
where ${\cal E}_{0}>0$ denotes the level of background numerical noises, and $\kappa>0$ is the Lyapunov exponent of chaotic system. Obviously, the smaller the background numerical noise ${\cal E}_{0}$, the larger the so-called “critical predictable time” $T_{c}$. This is the reason why CNS can give reliable simulations of chaotic systems and turbulent flows in a much larger interval of time $t\in[0,T_{c}]$ than DNS, because ${\cal E}_{0}$ of CNS is much smaller than that of DNS. From this viewpoint, we can regard DNS as a special case of CNS.
Finally, it should be emphasized that CNS results can provide us enlightenments to approach some mathematical truths such as the theorem 3 proved in this paper and the conjectures mentioned above and in [11]. Hopefully, CNS [12, 13, 7, 10, 9, 17, 20, 8, 18, 19, 11] could be used as a powerful tool to study turbulence.
Acknowledgements
This work is partly supported by National Natural Science Foundation of China (No. 12521002).
Data and materials availability
All data are available by sending requirement to the corresponding author.